# Make the Fastest Faster: Importance Mask for Interactive Volume Visualization using Reconstruction Neural Networks

### v1

@@ Abstract
Visualizing a large-scale volumetric dataset with high resolution is challenging due to the high computational time and space complexity. Recent deep-learning-based image inpainting methods significantly improve rendering latency by reconstructing a high-resolution image for visualization in constant time on GPU from a partially rendered image where only a small portion of pixels go through the expensive rendering pipeline. However, existing methods need to render every pixel of a predefined regular sampling pattern. In this work, we provide Importance Mask Learning (IML) and Synthesis (IMS) networks which are the first attempts to learn importance regions from the sampling pattern to further minimize the number of pixels to render by jointly considering the dataset, user’s view parameters, and the downstream reconstruction neural networks. Our solution is a unified framework to handle various image inpainting-based visualization methods through the proposed differentiable compaction/decompaction layers. Experiments show our method can further improve the overall rendering latency of state-of-the-art volume visualization methods using reconstruction neural network for free when rendering scientific volumetric datasets. Our method can also directly optimize the off-the-shelf pre-trained reconstruction neural networks without elongated retraining.
@@
The proposed IML and IMS Networks can further improve the rendering latency of already fast volume visualization methods using reconstruction neural networks, like the super-resolution EnhanceNet (a and b) and foveated rendering FoVolNet(c and d), with similar rendering quality. For each subfigure, the rightmost portion is the baseline using the traditional ray casting method, the leftmost portion is the interactive visualization method using reconstruction neural network, and the middle portion is the proposed method using our IML/S Net that empowers the respective reconstruction neural network.
@@
Introduction Interactive volume visualization techniques are crucial for enabling researchers across various disciplines to efficiently discover insightful features within scientific datasets from fields like medical imaging, geophysics, meteorology, materials science, and physical simulations. A visualization system that quickly responds to user interactions, adapting to data and viewing adjustments, can greatly enhance the efficiency and effectiveness of exploring complex scientific datasets. However, interactive volume visualization is challenging because of the following factors: First, the size of volumetric data generated nowadays is growing exponentially, necessitating an efficient storage and I/O system to facilitate large-scale data streaming between different computing units for generating comprehensive visualization results. Second, visualization techniques, like ray casting, ray tracing, global illumination, and high-order rendering, often demand significant computational resources to generate high-resolution, high-fidelity images.
@@
Various strategies have been proposed to address this performance issue in volumetric data visualization. For example, distributed volume visualization methods [35, 47] harness the parallel machines to divide the workload and then composite the partially rendered results into the final rendering image. Multi-resolution methods [12, 8, 9, 36] optimize the arrangement of data segments with different levels of detail (LOD) to improve the rendering latency while maintaining a high rendering quality. To improve the smoothness of volumetric data exploration, visualization systems also utilize caching and prefetching [7, 23] to reduce the data movement across the system memory hierarchy to decrease the input latency while rendering continuous frames along a user’s exploratory trajectory. The user behavior while exploring volumetric data can be learned [38, 13] to better assist the prefetching through the prediction of view parameters.
@@
Recent image synthesis techniques utilizing deep learning-based super-resolution in both spatial and temporal domains [15, 10, 19] are being adopted for visualizing large-scale volumetric datasets with responsive input latency. Given a partially rendered image from a view angle, where only a small percentage of the total pixels are rendered, the idea is to guess or interpolate the missing pixels through a trained neural network in constant time by leveraging the parallelization of GPU instead of going through the complex rendering algorithm to calculate the color blending for each pixel of the final image. The network that reconstructs the incomplete image to a full image is called a reconstruction neural network and we name such network as RecNN for the rest of the paper. Those methods rely on training an implicit neural representation [33, 34] from the original large-scale data considering complex rendering pipeline, and the trained representation is fast to inference for interactive visualization tasks. Volume visualization methods utilizing RecNNs give state-of-the-art performances in rendering latency [28, 3] when handling large-scale scientific datasets with complicated lighting effects.
@@
The bag of pixels to render for each partially rendered image is determined by a binary sampling pattern provided by the specific reconstruction neural network used (e.g., downsampling patterns for super-resolution RecNNs and foveal patterns for foveated rendering RecNNs). However, such a pattern is a predefined hyperparameter before training the network, and it does not adapt to any of the key visualization parameters like user’s views, transfer functions, and features of the underlying dataset. In this paper, we provide a method to further optimize the sampling pattern by finding which pixels contribute the most to the final reconstructed image (important pixels) and only rendering those pixels instead of all the pixels in the sampling pattern. The selection of important pixels from the sampling pattern is called the Importance Mask (IM). We proposed an Importance Mask Learning Network (IML Net) to learn the important pixels from the sampling pattern by jointly considering the dataset, user’s view parameters, and the downstream RecNN. The learned IM will be used to train an Importance Mask Synthesis Network (IMS Net), which will directly generate the importance mask from each new view during the user’s exploration. To the best of our knowledge, this is the first attempt at optimizing sampling patterns of visualization frameworks using RecNN. We use two state-of-the-art image inpainting RecNNs for volume visualization, EnhanceNet [28] for super-resolution and FoVolNet[3] for foveated rendering, to showcase how our method can help to further improve the rendering latency for free without sacrificing noticeable rendering quality. The main contributions of this work include:
@@
  • Importance Mask Learning Network(IML Net), a unified neural network to learn the importance mask from diverse sampling patterns by jointly considering the volumetric dataset, user’s view parameters, and the downstream reconstruction neural network.
@@
  • Importance Mask Synthesis Network(IMS Net), a regressor to directly predict the importance mask from the view parameters for fast generation of the partially rendered images.
@@
  • IML/S Net + RecNN, a visualization image synthesis network to further improve the already fast volume visualization methods using RecNN for free with similar rendering quality.
@@
  • Our networks can also be efficiently trained independently to the downstream RecNN to optimize the rendering latency for off-the-shelf pre-trained RecNN.
@@ 1.1 Implicit Neural Representation
Implicit Neural Representation (INR) refers to a method of representing data, such as images, 3D shapes, or other types of continuous signals, using a neural network. For handling large-scale volume visualization, compression methods utilizing implicit neural representation [22, 40] have been proposed to reduce network size and optimize I/O-intensive operations. For improving the rendering latency of the visualization system, Wu et al. [45] use hash encoding-based INR to speed up the interactive volume visualization with high reconstruction quality. Yariv et al. [46] improve geometry representation and reconstruction in neural volume rendering by modeling the volume density as an implicit function of the geometry. Weiss et al. [43] introduce fV-SRN as a novel extension of SRN (Scene Representation Networks) to achieve significantly accelerated reconstruction performance of volume rendering. Despite the use of powerful GPUs, the training time for INR remains significantly long when processing complex, large-scale scientific datasets [41, 45]. Both the proposed IML and IMS Nets were trained as INRs with learned properties of the dataset, view parameters, and the downstream RecNN.
@@ 1.2 Reconstruction Neural Network
A Reconstruction Neural Network (RecNN) in visualization refers to a type of neural network used to reconstruct or generate detailed representations of data from incomplete, noisy, or compressed input [20, 6]. The goal of RecNN is to recover or restore a high-quality, detailed version of an object or dataset from a simplified or partial form. The most commonly used type of RecNN is super-resolution neural networks that reconstruct high-res data from low-res data. The sampling pattern of super-resolution RecNN is a downsampling pattern. The EnhanceNet, proposed by Sajjadi et al. [28], demonstrates high reconstruction quality in generating volume visualization images [44]. Weiss et al. [42] propose a volumetric isosurface rendering with deep learning-based super-resolution. Tang et al. [41] propose STSR-INR which extends the super-resolution to both spatial and temporal domains. Another important RecNN is foveated rendering RecNN which does reconstruction from partial rendering near the region where the user is focusing the gaze. Together with the techniques in eye-tracking-related research [39, 49], foveated rendering provides a solution to balance the user experience and computational cost. The sampling pattern of foveated rendering RecNN is a foveal pattern. Kaplanyan et al. propose DeepFovea [17], the first adversarial neural network (GAN)-based foveated reconstruction method to speed up the rendering frame rate for gaming. Then Bauer et al. propose FoVolNet [3] to improve the foveated rendering in the volume visualization domain and achieve state-of-the-art rendering latency. In this work, we select two RecNNs, EnhanceNet and FoVolNet, as representations of super-resolution RecNN and foveated rendering RecNN to demonstrate how much rendering latency can be optimized using our method.
@@ 1.3 Image Synthesis
Deep learning-based image synthesis in volume visualization refers to the use of deep learning techniques to generate or enhance visual representations of 3D volumetric data, such as medical scans (MRI, CT), scientific simulations, or geological models [21]. Volumetric data can be very large, making real-time rendering computationally expensive. Generative models can synthesize results efficiently by learning the underlying structure of the data, reducing the need for extensive data processing and enabling faster generation of visual results. Berger et al. [4] use a GAN framework to directly generate the volume visualization image from view parameters and transfer functions. He et al. [11] provide a regressor to directly generate visualization images from simulation and visualization parameters for fast parameter space exploration. Jun et al. propose VCNet, a new deep-learning approach for volume completion by synthesizing missing subvolumes. Our IMS Net also utilizes image synthesis techniques to generate an importance mask directly from view parameters. Although training generative models, especially for large-scale or high-dimensional data, requires significant computational resources and large training datasets, the importance mask learned by the IML Net is a simple 2D binary array that can be easily learned by IMS Net.
@@ 2.1 Overview
For a general volume visualization system, the rendering latency (RL) is the time difference between the moment the user makes a data- or view-dependent operation and the time when the visualization image is done rendering. For interactive volume visualization using RecNN, as shown in Fig. 1, the rendering latency ($T_{RL}$) is the sum of two components which are the time of selective rendering (SR) for generating the partially rendered image (PR-Image), $T_{SR}$, and the inferencing time of the RecNN to generate a fully rendered image (FR-Image) from PR-Image, $T_{RecNN}$:
@@
$T_{RecNN}$ is determined by the specific RecNN architecture used in the reconstruction pipeline. We define the pixels covered by the sampling pattern as $S_{sp}$. For a sampling pattern with $I$ number of pixels, $T_{SR}$ is the sum of time used to render each pixel in the sampling pattern defined by the specific RecNN.
$T_{SR}=\sum_{k=0}^{I-1}T_{k}$ (2)
Once the RecNN is trained, the inferencing time $T_{RecNN}$ can be accelerated by GPU in constant time. The time used to render the PR-Image is the main overhead of the overall rendering latency. The time used to render the $k$th pixel, $T_{k}$, is determined by the complexity of the chosen rendering technique (ray casting [26], ray tracing [24], global illumination [50] etc.) and the interpolation method (linear or high-order [37]). Because of the high computational cost of pixel rendering, decreasing $I$ can significantly lower the rendering latency. However, $I$ is predefined by the sampling pattern which is a fixed hyperparameter for existing RecNNs. In order to further decrease the computational cost, a new set of important pixels $S_{im}$ needs to be selected from $S_{sp}$, and the rendering of $S_{im}$ can losslessly reconstruct the rendering of $S_{sp}$. We call the binary mask that filters out $S_{im}$ from $S_{sp}$ as Importance Mask(IM). For a chosen rendering function $R()$ and reconstruction function $f()$. The optimal $S_{im}$ can be derived from:
@@
where $\epsilon$ is a small predefined interval of error metric between the rendering of $S_{sp}$ and the reconstructed rendering from $S_{im}$. $\left|S_{im}\right|$ is the number of important pixels. Our objective is to select an optimal subset from the sampling pattern with the fewest possible pixels while ensuring that the rendering of this subset can accurately reconstruct the full sampling pattern with minimal reconstruction error. We name the percentage of pixels when using optimal $S_{im}$ as the optimal rendering percentage (ORP).
The volumetric datasets, mostly collected from scientific domains, show strong spatial correlations in their 3D domains. Reflected on the 2D image domain, such spatial correlation can also be observed. This provides opportunities to compress the complex pixel-level rendering into compact representations. Although there are numerous existing works to extract superpixels directly from 2D datasets [2], $S_{im}$ needs to be selected by not only considering the data itself but jointly considering the users’ input and downstream RecNN. Inspired by the saliency map [31] demonstrating which parts of the input data (e.g., pixels in an image) contribute most to the predictions of a neural network, we provide a method to automatically learn a selection of pixels that are the most important to reconstructing the rendering of sampling pattern. Instead of calculating the gradient of output with respect to the input, as the saliency map does, we utilize an autoencoder architecture to learn the IM as a latent feature through backpropagation from the training dataset. In this work, we propose two neural networks: Importance Mask Learning Network (IML Net) and Importance Mask Synthesis Network (IMS Net). IML Net will learn the importance mask from training datasets which are visualization images generated from perspective user view-dependent operations. Once the importance masks are learned, they will then be used as training datasets for the IMS Net which will learn the mapping from user’s view to the IM.

### v2

# Make the Fastest Faster: Importance Mask Synthesis for Interactive Volume Visualization using Reconstruction Neural Networks

@@ Abstract
Visualizing a large-scale volumetric dataset with high resolution is challenging due to the substantial computational time and space complexity. Recent deep learning-based image inpainting methods significantly improve rendering latency by reconstructing a high-resolution image for visualization in constant time on GPU from a partially rendered image where only a portion of pixels go through the expensive rendering pipeline. However, existing solutions need to render every pixel of either a predefined regular sampling pattern or an irregular sample pattern predicted from a low-resolution image rendering. Both methods require a significant amount of expensive pixel-level rendering. In this work, we provide Importance Mask Learning (IML) and Synthesis (IMS) networks, which are the first attempts to directly synthesize important regions of the regular sampling pattern from the user’s view parameters, to further minimize the number of pixels to render by jointly considering the dataset, user behavior, and the downstream reconstruction neural network. Our solution is a unified framework to handle various types of inpainting methods through the proposed differentiable compaction/decompaction layers. Experiments show our method can further improve the overall rendering latency of state-of-the-art volume visualization methods using reconstruction neural network for free when rendering scientific volumetric datasets. Our method can also directly optimize the off-the-shelf pre-trained reconstruction neural networks without elongated retraining.
@@
The proposed IML and IMS Networks can further reduce the rendering latency of already fast volume visualization methods using reconstruction neural networks, like the super-resolution EnhanceNet (a and b) and foveated rendering FoVolNet (c and d), with similar rendering quality. For each subfigure, the rightmost portion is the baseline using the traditional ray casting method, the leftmost portion is the interactive visualization method using reconstruction neural network, and the middle portion is the proposed method using our IML/S Net that enhances the performance of respective reconstruction neural network.
@@
Introduction Interactive volume visualization techniques are crucial for enabling researchers across various disciplines to efficiently discover insightful features within scientific datasets from fields like medical imaging, geophysics, meteorology, materials science, and physical simulations. A visualization system that quickly responds to user interactions, adapting to data and viewing adjustments, can greatly enhance the efficiency and effectiveness of exploring complex scientific datasets. However, interactive volume visualization is challenging because of the following factors: First, the size of volumetric data generated nowadays is growing exponentially, necessitating an efficient storage and I/O system to facilitate large-scale data streaming between different computing units for generating comprehensive visualization results. Second, visualization techniques, such as ray casting, ray tracing, global illumination, and high-order rendering, often require significant computational resources to generate high-resolution, high-fidelity images.
@@
Various strategies have been proposed to address this performance issue in large-scale volumetric data visualization. For example, distributed volume visualization methods [35, 48] harness the parallel machines to divide the workload and then composite the partially rendered results into the final rendering image. Multi-resolution methods [12, 8, 9, 36] optimize the arrangement of data segments with different levels of detail (LOD) to improve the rendering latency while maintaining a high rendering quality. To improve the smoothness of volumetric data exploration, visualization systems also utilize caching and prefetching [7, 23] to reduce out-of-core data movement across the system memory hierarchy, thereby decreasing input latency while rendering continuous frames along a user’s exploratory trajectory. The user behavior while exploring volumetric data can be learned [39, 13] to better assist the prefetching through the prediction of view parameters.
@@
Recent image synthesis techniques utilizing deep learning-based super-resolution in both spatial and temporal domains [15, 10, 19] are being adopted for visualizing large-scale volumetric datasets with responsive input latency. Given a partially rendered image from a view angle, where only a small percentage of the total pixels are rendered, the idea is to predict or interpolate the missing pixels through a trained neural network in constant time by leveraging the parallelization of the GPU instead of going through expensive rendering algorithms to calculate color blending for each pixel of the final image. The network that reconstructs an incomplete image into a full image is called a reconstruction neural network, and we refer to such a network as RecNN in the rest of this paper. Those methods rely on training an implicit neural representation [33, 34] using the original large-scale dataset and a complex rendering pipeline, and once trained, the representation enables fast inference, making it well-suited for interactive visualization tasks. Volume visualization methods based on RecNNs have achieved state-of-the-art performance in rendering latency [28, 3], particularly when processing large-scale scientific datasets with complex lighting effects.
@@
The bag of pixels to render for each partially rendered image is determined by a binary sampling pattern provided by the specific reconstruction neural network used (e.g., downsampling patterns for super-resolution RecNNs and foveal patterns for foveated rendering RecNNs). However, such a pattern is a predefined hyperparameter before training, and it does not adapt to any of the key visualization parameters like the user’s views, transfer functions, or features of the underlying dataset. Recently proposed Learning Adaptive Sampling (LAS) [45] tries to predict an arbitrary selection of important pixels, called the Importance Mask (IM), from a low-res rendering image. The ratio of the important pixels to the total rendering pixels is called the Important Pixel Ratio (IPR). Although LAS reduces the number of pixels to render for the downstream RecNN, it suffers from several key drawbacks that limit its practicality and lead to suboptimal performance gains: 1) Rendering the low-res input image still takes a relatively long time. 2) The IM can only be learned from the 2D space with the same resolution of the final rendering image, which can be very high for scientific visualization (normally $\geq 512\times 512$), resulting in a large number of samples to render for a specific IPR with a certain quality level. 3) LAS only supports the super-resolution type of RecNNs and needs a long time to retrain its weights. In this work, we provide a unified Importance Mask Learning Network (IML Net) to learn and distill the IM from a low-res intermediate embedding supporting various types of downstream RecNNs through the proposed compaction/decompaction layers. IML Net learns the IM directly from the sampling pattern by jointly considering the dataset, the user’s view parameters, and the downstream RecNN. We also proposed an Importance Mask Synthesis Network (IMS Net), which will directly synthesize the IM from novel views during the user’s exploration. Our neural rendering network (IML/S Net) combines the two networks with the downstream RecNN, achieving much faster rendering than LAS with similar rendering quality under the same IPR. We use two state-of-the-art image inpainting RecNNs for volume visualization, EnhanceNet [28] for super-resolution, and FoVolNet[3] for foveated rendering, to showcase how our method can further reduce the rendering latency for free without sacrificing noticeable rendering quality. We also perform a comprehensive comparison with LAS for both rendering quality and latency. The main contributions of this work include:
@@
  • Importance Mask Learning Network(IML Net), a unified neural network to learn the IM from low-res intermediate embedding derived from diverse sampling patterns by jointly considering the volumetric dataset, the user’s view parameters, and the downstream reconstruction neural network.
@@
  • Importance Mask Synthesis Network(IMS Net), a regressor designed to efficiently predict the importance mask directly from view parameters instead of computationally expensive input.
@@
  • IML/S Net + RecNN, a visualization neural rendering network to further improve the already fast volume visualization methods using RecNN for free with similar rendering quality.
@@
  • Our networks can also be efficiently trained independently of the downstream RecNN to optimize the rendering latency for off-the-shelf pre-trained RecNN.
@@ 1.1 Implicit Neural Representation
Implicit Neural Representation (INR) refers to a method of representing data, such as images, 3D shapes, or other types of continuous signals, using a neural network. For handling large-scale volume visualization, compression methods utilizing implicit neural representation [22, 41] have been proposed to reduce network size and optimize I/O-intensive operations. To improve the rendering latency of the visualization system, Wu et al. [46] utilize hash encoding-based INR to accelerate interactive volume visualization with high reconstruction quality. Yariv et al. [47] improve geometry representation and reconstruction in neural volume rendering by modeling the volume density as an implicit function of the geometry. Weiss et al. [44] introduce fV-SRN as a novel extension of SRN (Scene Representation Networks) to achieve significantly accelerated reconstruction performance of volume rendering. Despite the use of powerful GPUs, the training time for INR remains significantly long when processing complex, large-scale scientific datasets [42, 46]. Sun et al. [38] propose F-Hash to significantly speed up the convergence of training INR from large-scale time-varying scientific data through a feature-based multi-resolution Tesseract input encoding. Both the proposed IML and IMS Nets were trained as INRs with learned properties of the dataset, view parameters, and the downstream RecNN.
@@ 1.2 Reconstruction Neural Network
A Reconstruction Neural Network (RecNN) in visualization refers to a type of neural network used to reconstruct or generate detailed representations of data from incomplete, noisy, or compressed input [20, 6]. The goal of RecNN is to recover or restore a high-quality, detailed version of an object or dataset from a simplified or partial form. The most commonly used type of RecNN is super-resolution neural networks that reconstruct high-res data from low-res data. The sampling pattern of super-resolution RecNN is a downsampling pattern. The EnhanceNet, proposed by Sajjadi et al. [28], demonstrates high reconstruction quality in generating volume visualization images [45]. Weiss et al. [43] propose a volumetric isosurface rendering with deep learning-based super-resolution. Tang et al. [42] propose STSR-INR, which extends the super-resolution to both spatial and temporal domains. Another important RecNN is the foveated rendering RecNN, which does reconstruction from partial rendering near the region where the user is focusing their gaze. Together with the techniques in eye-tracking-related research [40, 50], foveated rendering provides a solution to balance the user experience and computational cost. The sampling pattern of foveated rendering RecNN is a foveal pattern. Kaplanyan et al. propose DeepFovea [17], the first foveated reconstruction method utilizing generative adversarial networks (GAN) to speed up the rendering frame rate for gaming. Then Bauer et al. propose FoVolNet [3] to improve the foveated rendering in the volume visualization domain and achieve state-of-the-art rendering latency. In this work, we select two RecNNs, EnhanceNet and FoVolNet, as representations of super-resolution RecNN and foveated rendering RecNN to demonstrate how much rendering latency can be further reduced using our method.
@@ 1.3 Image Synthesis
Deep learning-based image synthesis in volume visualization refers to the use of deep learning techniques to generate or enhance visual representations of 3D volumetric data, such as medical scans (MRI, CT), scientific simulations, or geological models [21]. Volumetric data can be very large, making real-time rendering computationally expensive. Generative models can efficiently synthesize rendering results by learning the underlying structure of the data, thereby reducing the need for expensive pixel-level computation and enabling faster generation of visual results. Berger et al. [4] utilize a GAN framework to directly generate the volume visualization image from view parameters and transfer functions. He et al. [11] provide a regressor to directly generate visualization images from simulation and visualization parameters for fast parameter space exploration. Jun et al. propose VCNet, a new deep-learning approach for volume completion by synthesizing missing subvolumes. Our IMS Net also utilizes image synthesis techniques to generate an importance mask directly from view parameters. Although training generative models, especially for large-scale or high-dimensional data, requires large training datasets and significant computational resources, the importance masks learned by the proposed IML Net are low-res 2D binary matrices that can be efficiently learned by the IMS Net.
@@ 2.1 Overview
For a general volume visualization system, the rendering latency (RL) is the time difference between the moment the user makes a data- or view-dependent operation and the time when the rendering of the visualization image is finished. For interactive volume visualization using RecNN, as shown in Fig.˜1, the rendering latency ($T_{RL}$) is the sum of two components, which are the time of Selective Rendering (SR) for generating the partially rendered image (PR-Image), $T_{SR}$, and the inferencing time of the RecNN to generate a fully rendered image (FR-Image) from PR-Image, $T_{RecNN}$:
@@
Once the RecNN is defined and trained, its inferencing time ($T_{RecNN}$) can be accelerated by GPU in constant time. We define the pixels covered by the sampling pattern as $S_{sp}$. For a specific sampling pattern, $T_{SR}$ is the sum of time used to render each pixel in the sampling pattern defined by the specific RecNN.
$T_{SR}=\sum_{k\in S_{sp}}T_{k}$ (2)
The time used to render the $k$th pixel, $T_{k}$, is determined by the complexity of the chosen rendering technique (ray casting [26], ray tracing [24], global illumination [51] etc.) and the interpolation method (linear or high-order [37]). The time used to render the PR-Image, $T_{SR}$, is the main overhead of the overall rendering latency. Because of the high computational cost of pixel rendering, decreasing the size of $S_{sp}$ can significantly reduce the rendering latency. However, the size of $S_{sp}$ is predefined by the sampling pattern, which is a fixed hyperparameter for existing RecNNs. In order to further decrease the computational cost, a new set of important pixels $S_{im}$ needs to be selected from $S_{sp}$, and the rendering of $S_{im}$ can losslessly reconstruct the rendering of $S_{sp}$. We refer to the binary mask that filters out $S_{im}$ from $S_{sp}$ as the Importance Mask (IM) for the remainder of the paper. For a chosen rendering function $R()$ and reconstruction function $f()$. The optimal $S_{im}$ can be derived from:
@@
where $\epsilon$ is a small predefined interval of error metric between the rendering of $S_{sp}$ and the reconstruction from the rendering of $S_{im}$. $\left|S_{im}\right|$ is the number of important pixels. Our objective is to select an optimal subset from the sampling pattern with the fewest possible pixels while ensuring that the rendering of this subset can accurately reconstruct the full sampling pattern with minimal reconstruction error. We name the percentage of pixels when using optimal $S_{im}$ as the Optimal Rendering Percentage (ORP).
The volumetric datasets, mostly collected from scientific domains, show strong spatial correlations in their 3D domains. Reflected on the 2D image domain, such spatial correlation can also be observed. This provides opportunities to compress the complex pixel-level rendering into compact representations. Although there are numerous existing works to extract superpixels directly from 2D datasets [2], $S_{im}$ needs to be selected by not only considering the data itself but also jointly considering the users’ input and downstream RecNN. Inspired by the saliency map [31] demonstrating which parts of the input data (e.g., pixels in an image) contribute most to the predictions of a neural network, we provide a method to automatically learn a selection of pixels that are the most important to reconstructing the rendering of the sampling pattern. Instead of calculating the gradient of output with respect to the input, as the saliency map does, we utilize an autoencoder architecture to learn the IM as a latent feature through backpropagation from the training dataset. In this work, we propose two neural networks: the Importance Mask Learning Network (IML Net) and the Importance Mask Synthesis Network (IMS Net). IML Net learns the importance mask from training datasets, which are visualization images generated from the respective user’s view-dependent operations. Once the importance masks are learned, they will then be used as a training dataset for the IMS Net, which will learn the mapping from the user’s view to the IM through novel view synthesis.
Figure 2: Demonstration of compaction and decompaction for super-resolution RecNN and foveated rendering RecNN.
Figure 3: Network architecture of IML Net. IML Net works on a compact 2D embedding domain (C-Image) with resolution of $n\times n$, which is much smaller than the full resolution output with a resolution of $m\times m$ ($m\gg n$).